PLEASE REPORT ANY OTHER PROBLEMS YOU NOTICE TO D. L. DePoy(firstname.lastname@example.org)!
This problem is caused by a slight mis-alignment of the f/7 camera optics with the collimated beam. The vignetting is due to the edge of the f/7 imaging aperture mask protruding into the beam. The mis-alignment can be corrected straightforwardly the next time the instrument is opened by adjusting the camera lenses and a folding flat.
Fixed by adjusting the first folding flat. The images
of the field was moved approximately 2 mm (March 1998)
This is a bug in the code that controls the display of the data. The bug occurred sometime between the initial check-out of the instrument in September and runs in November. We are currently working on fixing the code so that the latest image is displayed.
Installed new display code that displays the correct image (March 1998)
The currently installed H filter has a significant long-wavelength leak. The leak is roughly 10% at 3000-4500 nm. The high background at these wavelengths will overwhelm the natural background at H despite the thermal blocker installed as part of the detector mount. When using the H filter, be sure to put the PK-50 blocker in the beam as well. This will attenuate the long wavelength throughput of the instrument sufficiently to remove the effect of the leak. We are currently investigating replacing the H filter with one that is better blocked.
Sandwiched a piece of PK-50 long wavelength blocking material in
with the H filter in its cell (September 1997)
There have been reports that the Ks filter may not produce good images (May 1998)
The Ks filter clearly has problems. Either it is not flat, has
internal cracks, or ???